Circular No. 6077

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

X-RAY NOVA IN SCORPIUS R. M. Hjellming and M. Rupen, National Radio Astronomy Obser-

vatory, report the results of new Very Large Array radio observations of GRO J1655–40: “On Sept. 11.2 UT, a new addition to the radio counterpart of this x-ray transient appeared; this addition was seen as an added optically-thick component in the radio spectrum, with preliminary fluxes of 0.44, 0.41, 0.37, 0.3, and 0.25 Jy at 1.49, 4.9, 8.4, 14.9, and 22.5 GHz, respectively, for the total spectrum. It also appeared as a new source in 14.9- and 22.5-GHz images, with a location between the northeastern and southwestern components (IAUC 6073).”

S. Sazonov and R. Sunyaev, Space Research Institute, Moscow, on behalf of the Granat WATCH team, report: “GRO J1655–40 appeared bright in observations by the Granat WATCH all-sky monitor on Sept. 6– 12. Preliminary average flux estimates for Sept. 6–7, 8–9, and 9–12 are 490 ± 30, 560 ± 70, and 970 ± 30 mCrab, respectively, in the energy band 8–20 keV. Such a strong continuous growth of the flux is rather unusual for x-ray novae. A sharp decrease in the intensity was observed when the flux dropped from 1100 ± 130 mCrab on Sept. 8.6 UT to 290 ± 90 mCrab on Sept. 8.7.”

NOTICE CONCERNING POSITIONS OF NOVAE AND SUPERNOVAE

At the recent meetings of the IAU, Commissions 27 (Variable Stars) and 28 (on behalf of its Working Group on Supernovae) agreed that positions of novae and supernovae should be specified in 2000.0 coordinates (more specifically, the FK5/J2000.0 system) with effect from 1995 Jan. 1. The decision was endorsed by Commission 6 (Astronomical Telegrams), which oversees the publication of these Circulars.

V635 CASSIOPEIAE G. M. Hurst, The Astronomer, reports that G. Poyner (Birmingham,

England) has observed a brightening of V635 Cas, the optical counterpart of the x-ray transient/Be system 4U 0115+634 (cf. IAUC 4931, 4942). Selected visual observations: Aug. 13.96 UT, 15.3 (Poyner); 17.98, 15.1 (Poyner); 29.97, 15.3 (Poyner); Sept. 2.02, 14.7 (Poyner); 2.03, 14.3 (T. Vanmunster, Landen, Belgium); 5.94, 14.6 (Poyner); 6.98, 14.5 (Poyner); 8.87, 14.5 (Poyner); 8.87, 14.4 (J. Day, Leicester, England).

1994 September 13 Daniel W. E. Green